Alex2007:

If f(x)=3x+2 and g(x)=x2−x, find the value. f(4)=

carlosbigbain:

To find the value of f(4), we can substitute 4 into the function f(x) and solve. Let's do it step by step: First, let's write out the function f(x) as f(x) = 3x + 2. Now, substitute 4 for x in the function: f(4) = 3 * 4 + 2. Simplifying further: f(4) = 12 + 2. And the final answer is: f(4) = 14. So, the value of f(4) is 14.